OpenStudy (anonymous):

Anyone know much about Radius of convergence? I have having a problem understanding it.( How to find it really)

OpenStudy (anonymous):

It says it is the absolute value of (x-1)/R = absolute value of x

OpenStudy (anonymous):

Just no idea what that means. For example, with this equation.\[\sum_{0}^{\infty} nx^n/(x+4)\]

OpenStudy (anonymous):

is this \[\sum \frac{nx^n}{x+4}\]?

OpenStudy (anonymous):

hope it will help you @paintballcs http://www.sosmath.com/calculus/radcon/radcon02/radcon02.html